### v2

@@ 5.2.2 Real-world Problems
[…]
Table 2: Comparison of hypervolume values on real-world problems
In Table 2, SVH-MOL with the annealing schedule demonstrates superior performance, as A-SVH-{LS, TCH, STCH} consistently achieves higher HV values compared to the baseline methods. Across each combination of SVH and a scalarization method, our proposed SVH-MOL generally obtains better results than the corresponding baselines.
Our SVH-MOL successfully captures the complete Pareto front, ensuring both convergence and diversity. In Figure 3, we visualize the Pareto front on RE37 with three objectives. The results highlight the effectiveness of our framework in handling complex Pareto sets. The Pareto fronts generated by A-SVH-{LS, TCH, STCH} demonstrate better coverage compared to the baselines. Notably, in A-SVH-LS and A-SVH-TCH, the obtained Pareto sets cover most of the optimal solution space, unlike PHN-LS or PHN-TCH, which only capture a limited portion. This confirms the robustness and adaptability of our approach. Overall, SVH-MOL leverages the repulsive force term to promote diversity and preventing premature convergence.
Figure 3: Pareto fronts are generated on RE37.
5.3 Multi-task learning
In Multi-task learning, we compare our SVH-MOL performance with the baseline methods: PHN-LS, PHN-EPO, PHN-TCH, PHN-STCH, and PHN-HVI [20]. The multi-task experiments consist of two benchmarks: image classification (Multi-MNIST, Multi-Fashion, Multi-Fashion+ MNIST) [17], and Multi-Output Regression (SARCOS) [42]
Image Classification
We train a hypernetwork to generate the complete set of parameters for a target network that performs image classification across three datasets: Multi-MNIST, Multi-Fashion, and Multi-MNIST+Fashion. Each dataset defines a two-objective multi-task learning problem, where the goal is to find a set of optimal Pareto solutions. Specifically, the tasks involve classifying two overlapping handwritten digits in Multi-MNIST, two overlapping fashion items in Multi-Fashion, and one handwritten digit and one fashion item in Multi-MNIST+Fashion. We use both the LSTM [43] layer and the MLP layer in our core hypernetwork, and we employ a Multi-LeNet [6] for the target network. In the training process, we set the number of preference vectors $n=4$ and train for 500 epochs.
In table 3, our SVH-MOL achieves better performance than baseline approaches. In particular, our framework shows consistent and strong performance under various combinations of SVH and scalarizations, exceeding that of the baselines. While SVH-MOL clearly achieves a high HV (hypervolume) indicator, it exhibits poor diversity (lower spacing indicator, Figure 4). We contend that this limitation is reasonable, given the difficulty of simultaneously achieving both convergence and diversity in the set of optimal solutions. Our SVH-MOL helps identify a better Pareto front while maintaining diversity among the solutions.
In this experiment, we compare performance between vanilla SVH-MOL and annealed SVH-MOL. These results indicate that A-SVH-MOL achieves better optimization compared to SVH-LS and A-SVH-LS, as it achieves more convergence and diversity. However, in other cases, the improvements are not clearly observed. We hypothesize that this phenomenon may stem from the sensitivity of the annealing factor, which strongly influences the parameter generation of the hypernetwork, specially in multi-task learning scenarios where the number of parameters is very large. In our experiments, the hypernetwork tends to be highly sensitive and is easily affected by annealing, which may explain why SVH-MOL does not consistently outperform the other methods.
Method Multi-MNIST Multi-Fashion Fashion-MNIST
HV Spacing HV Spacing HV Spacing
PHN-LS 2.8056 0.0016 2.1292 0.0019 2.6915 0.0065
PHN-Chebyshev 2.8461 0.0081 2.1984 0.0159 2.7811 0.0324
PHN-SmoothChebyshev 2.8641 0.0028 2.1926 0.0082 2.7735 0.0211
PHN-EPO 2.7983 0.0059 2.1189 0.0114 2.6564 0.0139
[…]
SVH-LS 3.0147 7.3386e-5 2.1865 0.0002 2.9257 0.0002
A-SVH-LS 3.0263 0.0004 2.3421 0.0004 2.8963 0.0003
SVH-TCH 3.0952 0.0018 2.3820 0.0056 2.9291 0.0040
[…]
A-SVH-STCH 3.0641 0.0005 2.3915 0.0028 2.9081 0.0033
Table 3: Performance comparison across three datasets: Multi-MNIST, Multi-Fashion, Fashion-MNIST
Figure 4: Pareto fronts are generated by methods. We evaluate the result for 25 preference vectors.
Multi-Output Regression.
The SARCOS dataset is a widely used benchmark for multi-task learning (MTL) and regression, particularly in robotics and control. We evaluate our model on a 7-task setup, where each task corresponds to predicting the torque of one of the seven robot arm joints. Each input is a 21-dimensional vector comprising joint positions (7), velocities (7), and accelerations (7). The output is a 7-dimensional vector of joint torques. The dataset contains 44,484 training and 4,449 test examples. We reserve 10% of the training set for validation.
Our results, presented in Table 4, show that SVH-MOL achieves superior performance compared to all baseline methods. Notably, A-SVH-MOL effectively promotes diversity in the learned Pareto set. These results demonstrate that SVH-MOL performs remarkably well on high-dimensional multi-objective problems, achieving superior outcomes in both optimization quality and solution diversity. This indicates that our framework offers a promising approach for approximating the Pareto front in many-objective optimization tasks.
Method PHN-LS PHN-TCH PHN-STCH PHN-EPO PHN-HVI SVH-LS SVH-TCH SVH-STCH A-SVH-LS A-SVH-TCH A-SVH-STCH
[…]
Spacing 0.0025 0.0100 0.0021 0.0029 0.0376 0.0283 0.0754 0.0302 0.0448 0.1233 0.0381
Table 4: Performance comparison across different methods on the SARCOS dataset
5.4 Extended Results
5.4.1 Trade-off Between Convergence and Diversity
Figure 5: Multi-MNIST: Comparative Analysis of Alpha Parameter.
In our research, we leverage the second term in the gradient update to encourage repulsion among particles, thereby enhancing diversity within the solution set. This naturally leads to a bi-objective scenario, where the objectives of convergence and diversity are inherently conflicting. In our approach, we emphasize diversity, which presents a significant challenge when the algorithm struggles to converge to the optimal solutions. The coefficient $\alpha$ in Formula 7 is introduced as a hyperparameter to promote diversity. However, when varying $\alpha$ to study its impact on the shape and distribution of the Pareto set, the results are inconclusive. A similar issue has been observed in the study by [35].
In Figure 5, our proposed annealing SVH-MOL exhibits distinct behavior as the value of $\alpha$ is varied. A higher $\alpha$ encourages greater diversity on the Pareto front, as reflected in its broader shape; however, this brings about a reduced convergence quality. This result demonstrates that our annealing SVH-MOL is more suitable for controlling the impact of the alpha value in the optimizing process.
5.4.2 Effect of the Annealing Schedule
In Table 5, we compare our annealed SVH-MOL with a baseline that employs a vanilla annealing schedule, where the annealing process is repeated with a fixed period. We refer to this baseline as C-SVH, representing cyclical SVH. These results demonstrate the effectiveness and stability of our framework, as it performs better in almost all experiments. Furthermore, the use of an annealing strategy enhances the performance of SVH, with C-SVH consistently achieving higher results compared to SVH without the annealing strategy.
Dataset Cyclical SVH (C-SVH) Annealed SVH (A-SVH)
LS TCH STCH LS TCH STCH
[…]
RE34 0.9312 0.8683 0.8685 0.6657 0.9964 0.9986
[…]
RE37 0.7103 0.8159 0.8324 0.8235 0.8320 0.8383
Table 5: Comparison of annealed SVH-MOL and cyclical SVH (C-SVH)
6 Conclusion
@@ Acknowledgment
This research was funded by Vingroup Innovation Foundation (VINIF) under project code VinIF.2024.DA113. The authors also want to acknowledge the support of the Center for Environmental Intelligence at VinUniversity (project VUNI.CEI.FS_0007).

### v3

@@ 5.5.1. Effect of Diversity Parameter $\alpha$
Figure 7. Effect of kernel bandwidth ($h$) on Multi-MNIST performance. (Left) Table of final Hypervolume (HV) values, comparing various fixed bandwidths to the adaptive median heuristic. (Right) Convergence curves for each bandwidth setting. The median heuristic (black line) provides the most robust and highest-performing result.
5.5.2. Effect of Kernel Bandwidth $h$
Figure 7 presents a comprehensive analysis of kernel bandwidth selection, comparing fixed values with the adaptive median heuristic approach. The bandwidth parameter $h$ critically influences particle interactions during training, governing the trade-off between local exploitation and global exploration.
As shown in the table, the median heuristic bandwidth achieves the best hypervolume performance ($3.0942\pm 0.0106$). Analysis of fixed bandwidth values reveals a clear trend: very small bandwidths ($h=1e^{-6}$ to $h=1e^{-4}$) maintain competitive performance close to the median heuristic, with $h=1e^{-5}$ achieving $3.0902\pm 0.0017$. However, as bandwidth increases beyond $h=1e^{-3}$, performance begins to degrade, with $h=1e^{-2}$ dropping to $2.9275\pm 0.0040$ and $h=1e^{-1}$ further declining to $2.9879\pm 0.0509$. Interestingly, $h=1e^{0}$ shows a partial recovery to $3.0717\pm 0.0141$, though still underperforming the median heuristic. These results demonstrate that while small fixed bandwidths can achieve reasonable performance, the median heuristic provides superior and more robust results by adaptively adjusting to the particle distribution, avoiding the need for manual tuning and ensuring consistent convergence. The convergence curves (right panel) provide additional insights, showing that while some configurations achieve rapid initial progress, all experience optimization instability during early training phases. These results underscore the importance of adaptive bandwidth selection, as the median heuristic automatically adjusts to the particle distribution, avoiding the drawback of manually tuned fixed values.
5.5.3. Effect of Number of Particles
Table 8 presents the results of varying the number of particles in the training process and its influence on the hypervolume (HV) performance. As the number of particles increases, the computational burden of both the kernel function evaluation and the gradient update step grows substantially, leading to longer training times per epoch. Furthermore, we observe a decreasing trend in the hypervolume as the number of particles increases. In particular, while a moderate number of particles can enhance exploration and diversity, an excessively large set appears to dilute the model’s optimization effectiveness, resulting in lower HV values. These findings suggest that simply increasing the sample set does not necessarily yield better performance. Instead, there exists an optimal balance between computational efficiency and optimization quality. Notably, our SVH-PSL model achieves its best performance with a relatively small number of particles, demonstrating that it can efficiently approximate the Pareto front without relying on a large sample set.
Particles ($n$) HV Time/Epoch (s)
@@ 5.5.3. Effect of Number of Particles
8 3.0877 $\pm$ 0.0051 145.30
16 3.0808 $\pm$ 0.0051 447.23
32 2.9950 $\pm$ 0.0037 1603.27
Table 8. Ablation study on the number of particles ($n$) used during training on the Multi-MNIST dataset. The table shows the trade-off between performance (Hypervolume, HV) and computational cost (Time/Epoch in seconds).
6. Limitations and Future Work
While SVH-MOL demonstrates strong performance across various multi-objective problems, several limitations warrant discussion:
Computational Complexity
The pairwise kernel computations in SVGD scale quadratically with the number of particles ($O(n^{2})$), which can become prohibitive for large particle counts. While our hypernetwork approach mitigates this through parameter sharing, very large-scale applications may require approximate kernel methods or sampling strategies.
Hyperparameter Sensitivity
Our method requires careful tuning of several hyperparameters, particularly the diversity parameter $\alpha$ and annealing schedule parameters ($T_{0}$, $\tau$). Although we provide extensive sensitivity analysis, automatic hyperparameter optimization would enhance practical usability.
High-Dimensional Objective Spaces
While we demonstrate effectiveness on problems with up to 7 objectives, the ”curse of dimensionality” may affect performance in many-objective optimization (problems with ¿10 objectives). The repulsive force mechanism may become less effective in very high-dimensional spaces.
Theoretical Guarantees
Although empirically effective, theoretical convergence guarantees for our annealed SVGD approach in multi-objective settings remain an open question. Formal analysis of convergence properties would strengthen the methodological foundation.
Future Directions
Although our approach is effective in finding the convergence and diversity Pareto set, a fixed kernel selection is not universal. We suggest that exploring an adaptive kernel might be more suitable in future work. This adaptive kernel eliminates a set of hyperparameters, using learnable parameters instead in kernel calculation. Besides reducing computational costs, this approach automatically optimizes the trade-off between the driving and repulsive terms. Following our SVH-MOL framework, the kernel matrix is defined in the objective space, where an effective adaptive kernel can capture the geometric structure of the objective dimensions and promote moving to the optimal region better.
7. Conclusion
@@ Acknowledgment
This research was funded by Vingroup Innovation Foundation (VINIF) under project code VinIF.2024.DA113.